OpenStudy (anonymous):

If the sides of a square are lengthened by 8cm, the area becomes 256cm^2. find the length of a side of the original square. The length of a side of the original square is?

OpenStudy (bahrom7893):

(x+8)^2=256

OpenStudy (bahrom7893):

x+8=16 x=8

OpenStudy (anonymous):

the length of a side of the original square is 8

OpenStudy (mayliz):

(x+8)(x+8)=256 x=8 or x=-24 Length of any positive x=8

OpenStudy (anonymous):

so my answer is 24

OpenStudy (bahrom7893):

NO 8

OpenStudy (anonymous):

ok I amjust confused a square has 4 sides if each side was only eight then you addd eith to each side that gives us 16 for each side giving us a total of 64 then to the ^2 giving us 128cm

OpenStudy (bahrom7893):

no area of a square is s^2, where s is side. I said s=8 8+8=16 A=16^2=256

OpenStudy (anonymous):

ok I think I got it thank you for you help